I need help on this question...FIND THE WIDTH AND LENGTH OF EACH RECTANGLE.
The width is one half the length. The perimeter is 54 cm.

Answer by mananth(16949)

let length be 2x
width will be x
Perimeter = 2*(L+W)=54
2*(2x+x)=54
2*3x=54
6x=54
x=9 width
length = 18